# Tuning spin-density separation via finite-range interactions: Dimensionality-driven signatures in dynamic structure factors

### v1

@@ III EQUATION OF STATE IN $d$-DIMENSIONAL SYSTEM
$\Omega_{\text{g}}\left(\mu,\sqrt{n_{0}}\right)=\frac{1}{2}\sum_{\bm{k},\pm}\left[E_{\pm}+\frac{2}{\beta}\ln\left(1-e^{-\beta E_{\pm}}\right)\right].$ (22)
In what follows, our primary focus lies on the EOS of the model system at zero temperature, i.e. the ground state energy and quantum depletion.
@@ III.1 $d=3$
In this subsection III.1, we plan to investigate the EOS of the model system in 3D. To this end, we compute the ground state energy of the model system, $E_{\text{g}}=\Omega_{0}+\Omega_{\text{g}}+2L^{3}\mu n_{0}$. Then, the expression for the ground-state energy can be written as follows:
@@
$\frac{\Delta N}{N}=\frac{1}{(2\pi)^{2}}\left(\frac{2mg^{\left(0\right)}}{\hbar^{2}}\right)^{\frac{3}{2}}\left(g^{\left(0\right)}n\right)^{\frac{1}{2}}I\left(z_{\rho,\sigma}\right),$ (28)
where the scaling function $I\left(z\right)$ is defined as
$I\left(z_{\rho,\sigma}\right)=$displaystyle-- $\frac{\sqrt{2}\left(z_{\sigma}-1\right)\left(1-y\right)^{\frac{3}{2}}}{3\left(1+z_{\sigma}\right)^{2}}$ (29)
@@
$\frac{\Delta N}{N}=\frac{1}{(2\pi)^{2}}\left(\frac{2mg^{\left(0\right)}}{\hbar^{2}}\right)^{\frac{3}{2}}\left(g^{\left(0\right)}n\right)^{\frac{1}{2}}I^{(0)}\left(y\right),$ (30)
with the scaling function $I^{(0)}\left(y\right)$ reading
$I^{(0)}\left(y\right)=\frac{\sqrt{2}}{3}\left[\left(1-y\right)^{\frac{3}{2}}+\left(y+1\right)^{\frac{3}{2}}\right].$ (31)
@@ III.2 $d=1$
Figure 3: Dimensionless scaling function $h^{(\text{1D})}(z_{\rho},z_{\sigma})$ defined in Eq. (34), with $z_{\rho}\equiv 4mng_{\rho}^{(2)}/\hbar^{2}$ continuously scanned in $[0,0.7]$ and $z_{\sigma}\equiv 4mng_{\sigma}^{(2)}/\hbar^{2}$ varied over discrete values in $[0,0.4]$.
$\frac{E_{\text{g}}^{\left(\text{1D}\right)}}{L^{3}}$ $=$ $g^{\left(0\right)\left(\text{1D}\right)}n_{0}^{2}+g^{\left(0\right)\left(\text{1D}\right)}_{12}n_{0}^{2}$ (33)
@@
The outcome in Eq. (34) aligns excellently with the corresponding one in Ref. [47]. When the parameter $y^{\left(\text{1D}\right)}=0.2$, the corresponding result of Eq. (34) are presented in Fig. 3. At this point, we verify the accuracy of our results. In the limit where the finite-range interaction vanishes, the scaling function reduces to the following form
@@
Eqs. (24), (28), and (LABEL:GSE1) constitute a pivotal result of this study, delivering explicit analytical expressions for the LHY corrections to the ground-state energy and quantum depletion influenced by finite-range interactions in 3D and quasi-1D systems, thereby elucidating the impact of effective interaction strengths on the EOS and providing critical insights for precisely tuning the separation of the DSF with two degrees of freedom in Sec. IV.
@@ IV.1 Density and spin excitation
In subsection IV.1, our goal is to investigate how finite-range interactions can influence density and spin excitations in the system. For this purpose, we proceed to express the bosonic field $\psi_{i}$ in terms of number density $n_{i}$ and phase $\theta_{i}$ as $\psi_{i}=\sqrt{n_{i}}e^{i\theta_{i}}=\sqrt{n_{0}+\delta n_{i}}e^{i\theta_{i}}$, where $n_{0}$ is the equilibrium density and $\delta n_{i}$ denotes density fluctuations. To characterize spin-density separation, we transform to the basis $\delta n_{\rho(\sigma)}=(\delta n_{1}\pm\delta n_{2})/\sqrt{2}$ and $\theta_{\rho(\sigma)}=(\theta_{1}\pm\theta_{2})/\sqrt{2}$, corresponding to the density ($\rho$) and spin ($\sigma$) degrees of freedom. Within this framework, the Gaussian-order action in Eq. (3) in real time is reformulated for a $d$-dimensional system ($d=1,3$) as follows:
$S_{\text{g}}$ $=$ $\int d^{3}\bm{r}\int dt\sum_{\gamma=\rho,\sigma}\Bigg{[}\frac{n_{0}\left(\nabla_{r}\theta_{\gamma}\right)^{2}}{2m}+\delta n_{\gamma}\partial_{t}\theta_{\gamma}$ (36)
@@
with the coupling strength $g_{\rho\left(\sigma\right)}^{\left(0\right)}=g^{\left(0\right)}\left(1\pm y\right)$. The reduced mass $m_{\rho(\sigma)}=m/(1+z_{\rho(\sigma)})$ accounts for the effect of finite-range interactions. When $z_{\rho(\sigma)}=0$, the effective action reduces to that of Ref. [18]. We emphasize that in the Gaussian-order action in Eq. (36), finite-range interaction unlike other interaction types decouple exclusively into the kinetic term governing particle number density fluctuations, resulting in a corresponding effective mass correction.
@@
with $Q=\left(\bm{k},\omega\right)$ is $3+1$ vector denoting the momenta $\bm{k}$ and the frequency $\omega$, and $\bm{\Psi}\left(Q\right)=\left(\delta n_{\rho},\theta_{\rho},\delta n_{\sigma},\theta_{\sigma}\right)$. Furthermore, the above matrix $\bm{M}^{\prime}$ can be deduced into
@@
$M_{\rho\left(\sigma\right)}$ $=$ $\left(\begin{array}[]{cc}\frac{\hbar^{2}k^{2}}{8m_{\rho\left(\sigma\right)}n_{0}}+\frac{g_{\rho\left(\sigma\right)}^{\left(0\right)}}{2}&\frac{1}{2}i\omega_{\rho\left(\sigma\right)}\\ -\frac{1}{2}i\omega_{\rho\left(\sigma\right)}&\frac{\hbar^{2}k^{2}n_{0}}{2m}\end{array}\right),$ (41)
@@ IV.2.1 Density and spin DSF in 3D
In Eq. (44), the $\chi_{\rho\left(\sigma\right)}$, $v_{\rho\left(\sigma\right)}$ and $\Gamma$ are referred as to the compressibility, sound speed and damping rate respectively. In what follows, we plan to calculate these quantities one by one.
@@
In Eq. (50), the dimensionless coefficients are defined as $\alpha_{\rho\left(\sigma\right)}=v_{\rho\left(\sigma\right)}/v_{\rho,\text{3D}}^{\left(0\right)}$ and $\zeta_{\rho\left(\sigma\right)}=\chi_{\rho}^{\left(0\right)-1}/\chi_{\rho\left(\sigma\right)}^{-1}$, with the dependence of $\alpha_{\rho(\sigma)}$ on the interaction parameters $z_{\rho(\sigma)}$ depicted in Fig. 4(c). The momentum $k$ is rendered dimensionless, as outlined in Sec. II, where $k^{\prime}=\hbar k/\sqrt{gn_{0}m}$, with $\hbar/\sqrt{gn_{0}m}=1$. Consequently, the dimensionless frequency is given by $\omega^{\prime}=\omega/v_{\rho}^{\left(0\right)}k^{\prime}$. For $k^{\prime}=1.25$ and $y=0.3$, the density compressibility $\chi_{\text{$\rho$}}^{\left(0\right)}$ can be approximated as constant. The dimensionless DSF functions $J_{\rho(\sigma)}(\omega^{\prime})$ vary with frequency $\omega^{\prime}$ for different values of the interaction parameters $z_{\rho}$ and $z_{\sigma}$, as illustrated in Fig. 4(a). From Fig. 4(a), we observe that the DSF peaks for both density and spin modes shift toward lower frequencies with increasing finite-range interaction strength, indicating enhanced interaction effects. The density wave exhibits a broader peak due to stronger Beliaev damping, whereas the spin wave retains a narrower and sharper response, underscoring their distinct dynamic behaviors in 3D Bosonic mixtures.
@@ IV.2.2 Density and spin DSF in 1D
As a first step, the compressibility of $\chi_{\rho(\sigma),\text{1D}}$ is determined by $\chi^{-1}=\frac{1}{L}\frac{\partial^{2}E}{\partial n^{2}}$ at constant system length $L$. We need to calculate the contribution to the ground state energy of the density and spin excitations provided in Eq. (42), which are given as
@@ V CONCLUSION
This work establishes the EOS with finite-range interactions and elucidates their role in governing spin-density separation within Bosonic mixtures. In Sec. III, we compute the EOS for 3D and 1D systems, including LHY corrections to the ground-state energy and quantum depletion. Section IV visually demonstrates spin-density separation through DSFs, revealing that finite-range interactions—parametrized by $z_{\rho(\sigma)}$—enable continuous tuning of spectral peak positions in both spin ($\sigma$) and density ($\rho$) channels. This highlights a dimension-dependent inversion of peak shifts: in 1D systems, DSF peaks shift to higher frequencies with increasing interaction strength, whereas in 3D they shift to lower frequencies, underscoring dimensionality’s critical role in separation dynamics.
@@
  • 3D verification: Quantum depletion $\frac{N-N_{0}}{N}$ (Eq. (28)) is evaluated as $\approx 0.0028\times I(z_{\rho(\sigma)})$ using experimental parameters $n\approx 3\times 10^{14}\leavevmode\nobreak\ \text{cm}^{-3}$, $a\approx 2.7517\leavevmode\nobreak\ \text{nm}$ (Fig. 2(b)). The resulting $\mathcal{O}(10^{-3})$ depletion confirms Bogoliubov approximation validity [62, 63].
@@
  • Parameter validation: Effective ranges $r_{\mathrm{e}}$ tunable via dark-state control ($-1.4\times 10^{6}$ to $4.2\times 10^{6}\leavevmode\nobreak\ a_{0}$[66] yield $z_{\rho(\sigma)}\gg 1$, confirming our $[0,1]$ range covers physical regimes.
@@
In summary, this work explores the impact of finite-range interactions on the EOS and spin-density separation in Bosonic mixtures across 3D and quasi-1D regimes. At the Gaussian level, we establish that finite-range interactions decouple spin degrees of freedom, modifying the effective mass as analytically derived in Sec. IV. This non-local-interaction-driven decoupling manifests in distinct DSF peak shifts: to lower frequencies in 3D and higher frequencies in quasi-1D. By tuning the range parameters $r_{\mathrm{e}}$ and $r_{\mathrm{e}_{12}}$, the DSF peaks of density ($\rho$) and spin ($\sigma$) modes become continuously adjustable, enabling precise control of spin-density separation and significantly enriching dynamical behavior beyond contact-interaction limits. While we approximate finite-range strengths using the 1D coupling $g^{(2)}$ from Refs. [71, 60], deriving rigorously dimensionally-reduced interaction strengths for quasi-1D systems remains an open challenge. Incorporating second-order couplings $g^{(2)}$ from 3D reduction could enhance DSF prediction accuracy and experimental validation. By bridging theoretical insights with experimental accessibility through Bragg spectroscopy protocols, this work advances understanding of interaction-driven phenomena in quantum gases and establishes a foundation for exploring dimensionally-tuned exotic phases governed by finite-range interactions.
@@ ACKNOWLEDGMENTS
We thank Tao Yu, Ying Hu, and Biao Wu for stimulating discussions and useful help. This work was supported by the Zhejiang Provincial Natural Science Foundation (Grant No. LZ25A040004) and the National Natural Science Foundation of China (Grants No. 12074344).

### v2

@@ III GROUND-STATE ENERGY AND QUANTUM DEPLETION IN $d$-DIMENSIONAL SYSTEM
$\Omega_{\text{g}}\left(\mu,\sqrt{n_{0}}\right)=\frac{1}{2}\sum_{\bm{k},\pm}\left[E_{\pm}+\frac{2}{\beta}\ln\left(1-\mathrm{e}^{-\beta E_{\pm}}\right)\right].$ (22)
In what follows, our primary focus lies on the ground-state energy and quantum depletion of the model system at zero temperature.
@@ III.1 $d=3$
In this subsection III.1, we plan to investigate the ground-state energy and quantum depletion of the model system in 3D. To this end, we compute the ground-state energy of the model system, $E_{\text{g}}=\Omega_{0}+\Omega_{\text{g}}+2L^{3}\mu n_{0}$. Then, the expression for the ground-state energy can be written as follows:
@@
$\frac{\Delta N}{N}=\frac{1}{(2\pi)^{2}}\left(\frac{2mg^{\left(0\right)}}{\hbar^{2}}\right)^{\frac{3}{2}}\left(g^{\left(0\right)}n\right)^{\frac{1}{2}}I_{\text{QD}}\left(z_{\rho,\sigma}\right),$ (28)
where the scaling function $I_{\text{QD}}\left(z\right)$ is defined as
$I_{\text{QD}}\left(z_{\rho,\sigma}\right)=$ - $\frac{\sqrt{2}\left(z_{\sigma}-1\right)\left(1-y\right)^{\frac{3}{2}}}{3\left(1+z_{\sigma}\right)^{2}}$ (29)
@@
$\frac{\Delta N}{N}=\frac{1}{(2\pi)^{2}}\left(\frac{2mg^{\left(0\right)}}{\hbar^{2}}\right)^{\frac{3}{2}}\left(g^{\left(0\right)}n\right)^{\frac{1}{2}}I^{(0)}_{\text{QD}}\left(y\right),$ (30)
with the scaling function $I^{(0)}_{\text{QD}}\left(y\right)$ reading
$I^{(0)}_{\text{QD}}\left(y\right)=\frac{\sqrt{2}}{3}\left[\left(1-y\right)^{\frac{3}{2}}+\left(y+1\right)^{\frac{3}{2}}\right].$ (31)
@@ III.2 $d=1$
Figure 3: Dimensionless scaling function $h^{(\text{1D})}(z_{\rho},z_{\sigma})$ defined in Eq. (34), with $z_{\rho}\equiv 4mn_{0}g_{\rho}^{(2)}/\hbar^{2}$ continuously scanned in $[0,0.7]$ and $z_{\sigma}\equiv 4mn_{0}g_{\sigma}^{(2)}/\hbar^{2}$ varied over discrete values in $[0,0.4]$.
$\frac{E_{\text{g}}^{\left(\text{1D}\right)}}{L}$ $=$ $g^{\left(0\right)\left(\text{1D}\right)}n_{0}^{2}+g^{\left(0\right)\left(\text{1D}\right)}_{12}n_{0}^{2}$ (33)
@@
The outcome in Eq. (34) aligns excellently with the corresponding one in Ref. [49]. When the parameter $y^{\left(\text{1D}\right)}=0.3$, the corresponding result of Eq. (34) are presented in Fig. 3. At this point, we verify the accuracy of our results. In the limit where the finite-range interaction vanishes, the scaling function reduces to the following form
@@
Eqs. (24), (28), and (LABEL:GSE1) constitute a pivotal result of this study, delivering explicit analytical expressions for the LHY corrections to the ground-state energy and quantum depletion influenced by finite-range interactions in 3D and quasi-1D systems, thereby elucidating the impact of effective interaction strengths on the ground-state energy and quantum depletion and providing critical insights for precisely tuning the separation of the DSF with two degrees of freedom in Sec. IV.
Eqs. (24), (28), and (LABEL:GSE1) constitute a pivotal result of this study, delivering explicit analytical expressions for the LHY corrections to the ground-state energy and quantum depletion influenced by finite-range interactions in 3D and quasi-1D systems. The one-loop approximation captures leading-order quantum fluctuations, neglecting higher-order terms such as two-loop corrections, which are subdominant in the weakly interacting regime. This approach is controllable, as our results align with established limits (e.g., Ref. [49] for $z=0$), thereby elucidating the impact of effective interaction strengths on the ground-state energy and quantum depletion and providing critical insights for precisely tuning the separation of the DSF with two degrees of freedom in Sec. IV.
@@ IV.1 Density and spin excitation
In subsection IV.1, our goal is to investigate how finite-range interactions can influence density and spin excitations in the system. For this purpose, we proceed to express the bosonic field $\psi_{i}$ in terms of number density $n_{i}$ and phase $\theta_{i}$ as $\psi_{i}=\sqrt{n_{i}}\mathrm{e}^{\mathrm{i}\theta_{i}}=\sqrt{n_{0}+\delta n_{i}}\mathrm{e}^{\mathrm{i}\theta_{i}}$, where $n_{0}$ is the equilibrium density and $\delta n_{i}$ denotes density fluctuations. To characterize spin-density separation, we transform to the basis $\delta n_{\rho(\sigma)}=(\delta n_{1}\pm\delta n_{2})/\sqrt{2}$ and $\theta_{\rho(\sigma)}=(\theta_{1}\pm\theta_{2})/\sqrt{2}$, corresponding to the density ($\rho$) and spin ($\sigma$) degrees of freedom. Within this framework, the Gaussian-order action in Eq. (3) in real time is reformulated for a $d$-dimensional system ($d=1,3$) as follows:
$S_{\text{g}}$ $=$ $\int d^{d}\bm{r}\int dt\sum_{\gamma=\rho,\sigma}\Bigg[\frac{n_{0}\left(\nabla_{r}\theta_{\gamma}\right)^{2}}{2m}+\delta n_{\gamma}\partial_{t}\theta_{\gamma}$ (36)
@@
with the coupling strength $g_{\rho\left(\sigma\right)}^{\left(0\right)}=g^{\left(0\right)}\left(1\pm y\right)$. The reduced mass $m_{\rho(\sigma)}=m/(1+z_{\rho(\sigma)})$ accounts for the effect of finite-range interactions, with the specific derivation of the finite-range action term provided in Appendix C. When $z_{\rho(\sigma)}=0$, the effective action reduces to that of Ref. [19]. In Eq. (36), we emphasize that finite-range interactions, unlike other interaction types [19, 20], exclusively affect the kinetic term of density fluctuations in the Gaussian-order action, resulting in an effective mass correction. This decoupling ensures that finite-range effects modify only the number density fluctuation dynamics, without impacting the phase-related terms.
@@
with $Q=\left(\bm{k},\omega\right)$ is $d+1$ vector denoting the momenta $\bm{k}$ and the frequency $\omega$, and $\bm{\Psi}\left(Q\right)=\left(\delta n_{\rho},\theta_{\rho},\delta n_{\sigma},\theta_{\sigma}\right)$. Furthermore, the above matrix $\bm{M}^{\prime}$ can be deduced into
@@
$M_{\rho\left(\sigma\right)}$ $=$ $\left(\begin{array}[]{cc}\frac{\hbar^{2}k^{2}}{8m_{\rho\left(\sigma\right)}n_{0}}+\frac{g_{\rho\left(\sigma\right)}^{\left(0\right)}}{2}&\frac{1}{2}\mathrm{i}\omega_{\rho\left(\sigma\right)}\\ -\frac{1}{2}\mathrm{i}\omega_{\rho\left(\sigma\right)}&\frac{\hbar^{2}k^{2}n_{0}}{2m}\end{array}\right),$ (41)
@@ IV.2.1 Density and spin DSF in 3D
In Eq. (44), the $\chi_{\rho\left(\sigma\right)}$, $v_{\rho\left(\sigma\right)}$ and $\Gamma$ are referred to as the compressibility, sound speed and damping rate respectively. In what follows, we plan to calculate these quantities one by one.
@@
In Eq. (50), the dimensionless coefficients are defined as $\alpha_{\rho\left(\sigma\right)}=v_{\rho\left(\sigma\right)}/v_{\rho,\text{3D}}^{\left(0\right)}$ and $\zeta_{\rho\left(\sigma\right)}=\chi_{\rho}^{\left(0\right)-1}/\chi_{\rho\left(\sigma\right)}^{-1}$, with the dependence of $\alpha_{\rho(\sigma)}$ on the interaction parameters $z_{\rho(\sigma)}$ depicted in Fig. 4(c). The momentum $k$ is rendered dimensionless, as outlined in Sec. II, where $k^{\prime}=\hbar k/\sqrt{2gn_{0}m}$, with $\hbar/\sqrt{2gn_{0}m}=1$. Consequently, the dimensionless frequency is given by $\omega^{\prime}=\omega/v_{\rho}^{\left(0\right)}k^{\prime}$. For $k^{\prime}=1.25$ and $y=0.3$, the density compressibility $\chi_{\text{$\rho$}}^{\left(0\right)}$ can be approximated as constant. The dimensionless DSF functions $J_{\rho(\sigma)}(\omega^{\prime})$ vary with frequency $\omega^{\prime}$ for different values of the interaction parameters $z_{\rho}$ and $z_{\sigma}$, as illustrated in Fig. 4(a). From Fig. 4(a), we observe that the DSF peaks for both density and spin modes shift toward lower frequencies with increasing finite-range interaction strength, indicating enhanced interaction effects. The density wave exhibits a broader peak due to stronger Beliaev damping, whereas the spin wave retains a narrower and sharper response, underscoring their distinct dynamic behaviors in 3D Bosonic mixtures.
@@ IV.2.2 Density and spin DSF in 1D
As a first step, the compressibility of $\chi_{\rho(\sigma),\text{1D}}$ is determined by $\chi^{-1}=\frac{1}{L}\frac{\partial^{2}E}{\partial n^{2}}$ at constant system length $L$. We need to calculate the contribution to the ground-state energy of the density and spin excitations provided in Eq. (42), which are given as
@@ V CONCLUSION
This work establishes the ground-state energy and quantum depletion with finite-range interactions and elucidates their role in governing spin-density separation within Bosonic mixtures. In Sec. III, we compute the ground-state energy and quantum depletion for 3D and 1D systems, including LHY corrections to the ground-state energy and quantum depletion. Section IV visually demonstrates spin-density separation through DSFs, revealing that finite-range interactions—parametrized by $z_{\rho(\sigma)}$—enable continuous tuning of spectral peak positions in both spin ($\sigma$) and density ($\rho$) channels. This highlights a dimension-dependent inversion of peak shifts: in 1D systems, DSF peaks shift to higher frequencies with increasing interaction strength, whereas in 3D they shift to lower frequencies, underscoring dimensionality’s critical role in separation dynamics.
@@
  • 3D verification: Quantum depletion $\frac{N-N_{0}}{N}$ (Eq. (28)) is evaluated as $\approx 0.0028\times I(z_{\rho(\sigma)})$ using experimental parameters $n\approx 3\times 10^{14}\penalty 10000\ \text{cm}^{-3}$, $a\approx 2.7517\penalty 10000\ \text{nm}$ (Fig. 2(b)). The resulting $\mathcal{O}(10^{-3})$ depletion confirms Bogoliubov approximation validity [65, 66].
@@
  • Parameter validation: Effective ranges $r_{\mathrm{e}}$ tunable via dark-state control ($1.4\times 10^{6}$ to $4.2\times 10^{6}\penalty 10000\ a_{0}$, where $a_{0}$ is the Bohr radius) [69] yield $z_{\rho(\sigma)}\gg 1$, confirming our $[0,1]$ range covers physical regimes.
@@
In summary, this work explores the impact of finite-range interactions on the ground-state energy and quantum depletion and spin-density separation in Bosonic mixtures across 3D and quasi-1D regimes. At the Gaussian level, we establish that finite-range interactions decouple spin degrees of freedom, modifying the effective mass as analytically derived in Sec. IV. This non-local-interaction-driven decoupling manifests in distinct DSF peak shifts: to lower frequencies in 3D and higher frequencies in quasi-1D. By tuning the range parameters $r_{\mathrm{e}}$ and $r_{\mathrm{e}_{12}}$, the DSF peaks of density ($\rho$) and spin ($\sigma$) modes become continuously adjustable, enabling precise control of spin-density separation and significantly enriching dynamical behavior beyond contact-interaction limits. While we approximate finite-range strengths using the 1D coupling $g^{(2)}$ from Refs. [74, 63], deriving rigorously dimensionally-reduced interaction strengths for quasi-1D systems remains an open challenge. Incorporating second-order couplings $g^{(2)}$ from 3D reduction could enhance DSF prediction accuracy and experimental validation. By bridging theoretical insights with experimental accessibility through Bragg spectroscopy protocols, this work advances understanding of interaction-driven phenomena in quantum gases and establishes a foundation for exploring dimensionally-tuned exotic phases governed by finite-range interactions.
@@ ACKNOWLEDGMENTS
We thank Tao Yu, Ying Hu, and Biao Wu for stimulating discussions and useful help. This work was supported by the National Natural Science Foundation of China (Grants No. 12574301) and the Zhejiang Provincial Natural Science Foundation (Grant No. LZ25A040004).